Calculate the number of milliliters of 0.673 M Ba(OH)2 required to precipitate all of the Al3+ ions in 111 mL of 0.662 M Al(NO3)3 solution as Al(OH)3. The equation for the reaction is:


2Al(NO3)3(aq) + 3Ba(OH)2(aq) 2Al(OH)3(s) + 3Ba(NO3)2(aq)


1
Expert's answer
2021-10-07T02:19:01-0400

We use the balanced equation (it was correctly balanced as it was presented) to find the volume of Ba(OH)2 required by using conversion factors:


"111\text{ mL sol. }Al(NO_3)_3 \Big(\frac{0.662\text{ mol }Al(NO_3)_3)}{\text{1000 mL sol. }Al(NO_3)_3} \Big) \Big( \frac{3\text{ mol }Ba(OH)_2}{2\text{ mol }Al(NO_3)_3} \Big) \Big( \frac{\text{1000 mL sol. }Ba(OH)_2}{0.673\text{ mol }Ba(OH)_2} \Big) 
\\ \text{ }
\\ = \text{163.778 mL} \approxeq \text{164 mL sol. }Ba(OH)_2 \text{ 0.662 M}"


In conclusion, we will need 164 mL of the Ba(OH)2 solution to precipitate all the Al3+ ions as Al(OH)3.

Reference

  • Chang, R., & Goldsby, K. A. (2010). Chemistry. Chemistry, 10th ed.; McGraw-Hill Education: New York, NY, USA.